The time period of a simple pendulum of length l is measured in an elevator descending with acc. =g/3

Grade:12

1 Answers

j.vamsee gowtham
14 years ago

hai yaarrr iam vamsee.consider the case when lift is moving with uniform velocity downwards after some time it starts descending with an acc a=g/3.so we can write a1=g-g/3.t=2.22/7.squqre root ofl/a1.thereforeT=2.22/7.SUARE ROOT OFI/G-G/3. 

IN THE OTHER CASE WHEN LIFT MOVING UNIFORMLY IN UPWARD DIRECTION AFTER COME TIME IT DESENDS WITH AN ACCELERATION G/3.THEN WE CAN WRITET=2.22/7SUARE ROOT OF L/G+G/3.......

IAM SORRY IF ANY MISTAKES ARE FOUND I TRIED MY LEVEL BEST YAAAR